Solve -7(11-4w)-15w=-27-(50w-13w)
Answer by dfrazzetto(283)

Solve -7(11-4w)-15w=-27-(50w-13w)
First multiply everything out:
-77 + 28w - 15w = -27 - 50w + 13w
-77 + 13w = -27 - 37w
add 77 and 37w to both sides...
50w = 50
w=1
CHECK, plug w=1 into original equation:
-7(11-4)-15=-27-(50-13)
-7(7)-15=-27 - 37
-49-15 = -64
-64=-64 √